Problem 1.14 Question Help Charles Lackey operates a bakery in Idaho Falls, Idaho. Because of its excellent product and excellent location, demand has increased by 25% in the last year. On far too many occasions customers have not been able to purchase the bread of their choice. Because of the size of the store, no new ovens can be added. At a staff meeting, one employee suggested ways to load the ovens differently so that more loaves of bread can be baked at one time. This new process will require that the ovens be loaded by hand, requiring additional manpower. This is the only production change that will be made in order to meet the increased demand. The bakery currently makes 1,600 loaves per month. Employees are paid $8 per hour. In addition to the labor cost, Charles also has a constant utility cost per month of $800 and a per loaf ingredient cost of $0.40. Current multifactor productivity for 640 work hours per month 리.256 loaves/dollar (round your response to three decimal places).cant figure out how to compute the input costs (labor costs + ingredient + Utility)?
Answer #1

Present Output = 1,600 loaves

Present Input ($) = 640 x $8 (labor hrs.) + $800 (utility) + 1600 x $0.4 (ingredients) = $6,560

So, multifactor productivity = 1600 / 6560 = 0.244 loaves per dollar